OpenStudy (anonymous):

Slove the following system of equations: x-y=13 2x+y=2 what method would you use?

OpenStudy (ankit042):

add the two equation first to solve for x, then put the value of x in any of the equation

OpenStudy (anonymous):

so it would be (5,-8)?

OpenStudy (johnweldon1993):

Here I would use the elimination method....you can add these 2 equations and eliminate the variable 'y' x-y=13 +(2x+y=2) ----------- 3x = 15 and that solves for 'x' and then use that to solve for 'y'...

OpenStudy (johnweldon1993):

and yes that would be correct (5 , -8) is what I get too

OpenStudy (ankit042):

Yes Great job @LeahL1111

OpenStudy (anonymous):

Thank you so much @johnweldon1993 and @ankit042 :)))

OpenStudy (johnweldon1993):

Anytime :)